The percentage of a​ country's population 25 years and older with at least a high school diploma is shown in the table for selected years. Let
x=0
represent​ 1950,
x=10
represent​ 1960, and so on.
​(a) Use the data for 1950 and 2000 to find a linear function f that models the data.

​(b) Use the function from part​ (a) to approximate the percentage of the​ country's population 25 years and older who were at least high school graduates in 1995.
High School Graduates


Year
Percent


1950
34.7


1960
40.7


1970
54.4

1980
66.6
1990
77.9
2000
87.3
2005
88.4


(a) A linear function that models the data is f(x)=

Answer by Boreal(15235)

Using 1950 and 2000, 34.7 and 87.3, one has 50 years (x-axis) with a change in per cent of 52.6
This is a slope of 1.052.
Letting 1950=0 and 2000=50, a linear equation with those two data points is
y-y1=m(x-x1) m slope, (x1, y1) a point, and can use (0, 34.7)
y-34.7=1.052*(x)
y=1.052x+34.7, where x is the number of years after 1950.
For 1995, x=45
y=45*1.052+34.7=82.04% or 82.0%